OpenStudy (anonymous):

how do you find f(x) = 4 + 5x and g(x) = 2x – 1find f(f(x)) ?

OpenStudy (anonymous):

f(f(x))=f(4+5x)=4+5(4+5x)....basically substituting from the inside till you go outside ....